OpenStudy (anonymous):

Find the value of (0) when f '(x)=sin2t, f(π/2)=3

OpenStudy (anonymous):

the value of 0?

OpenStudy (anonymous):

Find the value of f(0) when f '(x)=sin2t, f(π/2)=3

OpenStudy (amistre64):

big-O lol

OpenStudy (amistre64):

antiderive sin(2t), or is that sin^2(t) ?

OpenStudy (anonymous):

sin(2t)

OpenStudy (amistre64):

sin(2t) we should know comes from at least a cos(2t) so what are we missing?

OpenStudy (anonymous):

\[F(x)=-\frac{1}{2}\cos(2x)+C\] and \[ f(\frac{\pi}{2})=-\frac{1}{2}\cos(\frac{\pi}{2})+C=3\]

OpenStudy (anonymous):

solve for C, then plug in 0 and see what you get

OpenStudy (amistre64):

Dt cos(2t) = -2 sin(2t)\[\] Dt -cos(2t) = 2 sin(2t)\[\] so we need to place a catch for that 2 that pops out\[\] -cos(2t)/2 seems to work; +C of course

OpenStudy (anonymous):

satellite did you use the u-substitution

OpenStudy (amistre64):

there is no magic to a "u-sub" .... its just a device to see the work easier

OpenStudy (anonymous):

so how did he or she come up with -1/2cos2x?

OpenStudy (amistre64):

its called; this is a simple enough antiderivative to be able to predict where it came from

OpenStudy (amistre64):

there is no "u-sub" magic :)

OpenStudy (amistre64):

if you want, you can work it all out in a u-sub

OpenStudy (amistre64):

antiD sin(2t) dt; u = 2t; du = 2 dt; du/2 = dt\[\] antiD sin(u) du/2 = -cos(u)/2; -cos(2t)/2

OpenStudy (amistre64):

but if you can see where the function came from to begin with and understand the mechanics that got it there; the rest of that is superflous